OpenStudy (anonymous):

\[xy'+y=y ^{2},y(1)=-1\]

OpenStudy (anonymous):

i reached \[2\tanh ^{-1}(1-2y)=\ln x+c\]

OpenStudy (anonymous):

any kind soul can guide me?

OpenStudy (turingtest):

Let me call on people that may be able to answer this: @across @lalaly @JamesJ @ash2326 it's just beyond me, sorry look forward to seeing the solution :)

OpenStudy (anonymous):

owh kk ty. actually i dont know what is tanh also. i reached the tanh based on the integration table ._. lol

OpenStudy (experimentx):

looks like Bernoulli's eqn

OpenStudy (anonymous):

ou im using separable method. anyway, can explain how \[2\tanh ^{-1}(1-2y)=\ln x+c\]becomes\[\ln(1-y) -\ln y=\ln x+c\] basically its the LHS tat i dont understand

OpenStudy (anonymous):

ou.. ._. never knew there's such thing. TQ! =)

OpenStudy (anonymous):

er, but if z=1-2y, why after becoming ln its 1-y

OpenStudy (experimentx):

should be 1/2*2(ln(1+1-2y) - ln(2y)) = ln2(1-y) - ln2y = ln 2 + ln(1-y) - ln 2 - ln y

OpenStudy (turingtest):

\[\tanh ^{-1}z=\frac12\ln(1+z)-\ln(1-(1-2y))\]\[\tanh ^{-1}z=\frac12[\ln(1+1-2y)-\ln(1-z)]=\ln2+\frac12\ln(1-y)-\frac12\ln y\]yep, experimentX beat me to it again ln2 is a constant which can be absorbed into C is the point...

OpenStudy (experimentx):

you missed ln 2 at last .. lol

OpenStudy (anonymous):

owh. now i get it... ._. TQVM =)

OpenStudy (turingtest):

d'oh, I was racing you too much :/ so they just cancel anyway...